Find a 2 digit number whose tens digit is 3 less than the units digit. The orginal number is 5 more than 4 times the sum of the digit
Answer by stanbon(48558)

Find a 2 digit number whose tens digit is 3 less than the units digit. The orginal number is 5 more than 4 times the sum of the digit
---------------
Let the number be 10t+u where t is the tens digit and u is the units digit.
------------
EQUATIONS:
t = u-3
10t+u = 4(t+u)+5
-------------
Rearranging:
t = u -3
6t-3u = 5
---------
Substitute to solve for "u":
6(u-3)-3u=5
6u-18-3u=5
3u=23
u=23/3
----------
Comment:
The condtions are not consistent; the value of u should be a whole number.
-----------
Cheers,
Stan H.